Find the basis and dimension of the vectors are (1,-3,1) ( 2,-6,2) and (3,-9,3) .


1
Expert's answer
2021-06-07T12:14:46-0400

Consider the augmented matrix


"\begin{bmatrix}
 1 & 2 & 3 & 0 \\
 -3 & -6 & -9 & 0 \\
 1 & 2 & 3 & 0
\end{bmatrix}"

"R_2=R_2+3R_1"


"\begin{bmatrix}
 1 & 2 & 3 & 0 \\
 0 & 0 & 0 & 0 \\
 1 & 2 & 3 & 0
\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & 2 & 3 & 0 \\
 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0
\end{bmatrix}"

Three vectors are collinear.

The basis vector is "(1, -3, 1)," for example. The dimension of basis is 1.